Journal of the Association for Computing Machinery, Vol. 42, No. 5, September 1995 (To appear)

Recoverable Sequence Transmission Protocols

EWAN D. TEMPERO

AND

RICHARD E. LADNER

Abstract. We consider the sequence transmission problem, that is, the problem of transmitting an infinite sequence of messages x_{1}x_{2}x_{3} … over a channel that can both lose and reorder packets. We define performance measures, ideal transmission cost and recovery cost, for protocols that solve the sequence transmission problem. Ideal transmission cost measures the number of packets needed to deliver x_{n} when the channel is behaving ideally and recovery cost measures how long it takes, in terms of number of messages delivered, for the ideal transmission cost to take hold once the channel begins behaving ideally. We also define lookahead, which measures the number of messages the sender can be ahead of the receiver in the protocol. We show that any protocol with constant recovery cost and lookahead requires linear ideal transmission cost. We describe a protocol, P_{lin} that has ideal transmission cost 2n, recovery cost 1, and lookahead 0.

Copyright 1995 by ACM, Inc. Permission to make digital or hard copies of part or all of this work for personal or classroom use is granted without fee provided that copies are not made or distributed for profit or direct commercial advantage and that copies bear this notice and the full citation on the first page. Copyrights for components of this work owned by others than ACM must be honored. Abstracting with credit is permitted. To copy otherwise, to republish, to post on servers, or to redistribute to lists, requires prior specific permission and/or a fee. Request permissions from Publications Dept, ACM Inc., fax +1 (212) 869-0481, or permissions@acm.org.